### v2

# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

Abstract.
We propose a novel algorithm for calculating the ground-state energy of quantum many-body systems by combining auxiliary-field quantum Monte Carlo (AFQMC) with tensor-train sketching. In AFQMC, having a good trial wavefunction to guide the random walk is crucial for avoiding sign problems. Typically, this trial wavefunction is fixed throughout the simulation. Our proposed method iterates between determining a new trial wavefunction in the form of a tensor train, derived from the current walkers, and using this updated trial wavefunction to anchor the next phase of AFQMC. Numerical results demonstrate that our algorithm is highly accurate for large spin systems, achieving a relative error of $10^{-5}$ in estimating ground-state energies. Additionally, the overlap between our estimated trial wavefunction and the ground-state wavefunction achieves a high-fidelity. We provide a convergence proof, highlighting how an effective trial wavefunction can reduce the variance in the AFQMC energy estimate.
1. Introduction
The quantum many-body problem appears in a wide range of fields, including condensed matter physics, high energy and nuclear physics, quantum chemistry, and material science. One of the most challenging parts of this problem is that the computational cost grows exponentially with the size of the system. Quantum Monte Carlo (QMC) [1, 2, 3, 4] is a class of algorithms that can efficiently deal with high-dimensional problems by reducing the computational cost to a polynomial scale with the system size. It guarantees that in expectation, one can obtain the exact ground-state energy. However, with a finite sample size, QMC generally suffers from sign problem [5, 6], meaning the variance grows exponentially in time. To fix this issue, the constrained-path auxiliary-field quantum Monte Carlo (cp-AFQMC) [7] and phaseless auxiliary-field quantum Monte Carlo (ph-AFQMC) [8] were developed to avoid the sign problem, and have been successfully applied to both lattice models [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and realistic materials [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. In cp-AFQMC, the wavefunction is represented as a sum of an ensemble of statistical independent random walkers. The sign problem is avoided by introducing a trial wavefunction to bias the random walkers such that they point in the same direction as the trial wavefunction, with a price of a potential systematic bias. This bias depends on the quality of the trial wavefunction [29]. Ideally, the bias will be removed if the trial wavefunction is exactly the ground-state wavefunction of the system [30]. Based on this result, self-consistent approaches for successively improving the trial wavefunction have been developed [31, 32, 33, 34] to reduce the systematic error.
On the other hand, a different type of approach towards the quantum many-body problem is to directly solve for the wavefunction through approximating it by some low-complexity ansatz. One of the most popular ansatz is the matrix product state (MPS) [35, 36], also known as the tensor train (TT) [37]. A huge advantage of TT representation is that its storage complexity and computational complexity can be made near linear (instead of exponential) in the dimension $d$. However, for complicated systems, the approximation error can be large, which leads to inaccurate energy estimates.
Recently, an algorithm that directly combines AFQMC and TT representation has been proposed by one of the authors [38]. In this method, one projects the random walkers into a TT by a TT-sketching technique [39] at every step. It presents a potentially cheaper strategy to update a TT, using AFQMC to simulate a matrix-vector multiplication. While this method avoids the expensive compression of the Hamiltonian operator, it still suffers from the limitation that sometimes one cannot approximate the ground-state wavefunction as a low-rank TT.
In this paper, we provide improvements to AFQMC and tensor methods by combining the best of both worlds. A TT representation of the wavefunction is estimated periodically using TT-sketching. Then the estimated tensor-train wavefunction is used as the trial wavefunction to “anchor” the next episode of AFQMC. Different from a previously proposed method that uses a fixed TT trial wavefunction [40], we alternate between estimating a new trial wavefunction, and running the next episode of AFQMC. While this can potentially remove the energy bias by successively improving the trial wavefunction, we show that this strategy also significantly reduces the energy variance as well. On the other hand, our approach in estimating the tensor-train wavefunction is made cheap by the fact that it does not have to be highly accurate, but just have to be sufficiently accurate to guide the walkers. This is unlike other tensor-train approaches, including [38], where it is imperative to obtain an accurate approximation to the ground state since the tensor-train is used to obtain to final energy estimate. We also motivate our procedure by theoretically showing that an improved trial wavefunction can significantly lower the variance of AFQMC’s energy.
The rest of this paper is organized as follows. In Section 2, we introduce the procedure of cp-AFQMC for calculating the ground state of spin models, and also some preliminaries for TT and TT-sketching. Our proposed algorithm is then presented in Section 3. In Section 4, we provide a theoretical analysis for the convergence of cp-AFQMC, showing that the procedure can give an accurate energy estimate, despite the fact that the wavefunction itself has a large variance. The results of numerical experiments are shown in Section 5.
2. Preliminary
2.1. cp-AFQMC
In this section, we illustrate the application of cp-AFQMC to determine the ground-state energy of a spin model. The ground-state problem is stated as
(1) $\Psi_{0}=\underset{\Psi\in\mathbb{R}^{2^{d}}}{\arg\min}\ \langle\Psi,H\Psi\rangle,\text{ s.t. }\left\|\Psi\right\|_{2}^{2}=1,$
where $d$ is the number of spins. The ground-state wavefunction $\Psi_{0}$ can be obtained by an imaginary-time evolution, i.e.
(2) $\Psi_{0}\propto\lim_{\tau\to\infty}e^{-\tau H}\Psi_{\mathrm{tr}}.$
where $\Psi_{0}$ denotes the solution to (1), and $\Psi_{\mathrm{tr}}$ is some trial wavefunction picked in prior satisfying $\braket{\Psi_{\mathrm{tr}},\Psi_{0}}\neq 0$. Numerically, the limit in (2) can be calculated iteratively by
$\Psi^{(n+1)}=e^{-\Delta\tau H}\Psi^{(n)}.$
Applying $e^{-\Delta\tau H}$ to $\Psi^{(n)}$ amounts to applying a $2^{d}\times 2^{d}$ matrix to a vector of size $2^{d}$, which suffers from the curse-of-dimensionality. To solve this issue, QMC performs $e^{-\Delta\tau H}\Psi^{(n)}$ with a Monte Carlo method. In particular, cp-AFQMC represents the wavefunction as the following:
(3) $\Psi^{(n)}\approx\widehat{\Psi}^{(n)}=\sum_{k}{\phi}_{k}^{(n)},$
where each ${\phi}_{k}^{(n)}$ is a meanfield or separable state, and $e^{-\Delta\tau H}$ is decomposed as
(4) $e^{-\Delta\tau H}=\sum_{\vec{x}}P(\vec{x})B(\vec{x})=\mathbb{E}_{\vec{x}\sim P}B(\vec{x}),$
where $B(\vec{x}){\phi}_{k}^{(n)}$ can be computed cheaply, and it remains a meanfield or separable state. Then, one step of imaginary-time evolution can be approximated by
$e^{-\Delta\tau H}\widehat{\Psi}^{(n)}=\sum_{k}e^{-\Delta\tau H}{\phi}_{k}^{(n)}=\sum_{k}\sum_{\vec{x}}P(\vec{x})B(\vec{x}){\phi}_{k}^{(n)}\approx\sum_{k}{\phi}_{k}^{(n+1)}=:\widehat{\Psi}^{(n+1)},$
where
$\phi^{(n+1)}_{k}=B(\vec{x}^{(n)}_{k}){\phi}_{k}^{(n)},\quad\vec{x}^{(n)}_{k}\sim P.$
While the idea is straightforward, cp-AFQMC further incorporates importance sampling when sampling $\vec{x}_{k}^{(n)}$ and also restricting the path of the random walk in order to reduce the variance of estimating certain quantities using the random walkers. The starting point is a representation of the walkers of the following form
(5) $\widehat{\Psi}^{(n)}=\sum_{k}w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{O_{\mathrm{tr}}(\phi^{(n)}_{k})},$
where,
$O_{\mathrm{tr}}(\phi_{k}^{(n)})=\max\left\{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}},0\right\},$
and the weight $w_{k}^{(n)}$ is either positive, or
$w^{(n)}_{k}=0\ \text{if}\ O_{\mathrm{tr}}(\phi_{k}^{(n)})=0.$
With this representation, in cp-AFQMC it is possible to zero out some random walkers when $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}\leq 0$, hence constraining the path of AFQMC.
To perform imaginary-time evolution, we incorporate a new probability distribution function $\widetilde{P}$, defined for a walker $\phi_{k}^{(n)}$ with positive weight $w_{k}^{(n)}>0$ as
(6) $\widetilde{P}_{k}^{(n)}(\vec{x})=\frac{1}{\mathcal{N}_{k}^{(n)}}\frac{O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}P(\vec{x}),$
where $\mathcal{N}_{k}^{(n)}$ is the normalization constant. This new probability distribution function favors those walkers that have large overlap with the trial wavefunction.
A remark here is that if an auxiliary field $\vec{x}$ leads to non-positive overlap with the trial function, so that $O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})=0$, then the probability is set to be $\widetilde{P}_{k}^{(n)}(\vec{x})=0$, meaning that this $\vec{x}$ will not be sampled anymore. In this way, we only sample from those $\vec{x}$ such that the walker at the next step $B(\vec{x})\phi_{k}^{(n)}$ has a positive overlap with the trial function, and thus a positive weight. Therefore, the new walker $B(\vec{x})\phi_{k}^{(n)}$ is guaranteed to have a positive weight, as long as $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for at least one $\vec{x}$.
With the new probability distribution function (6), the operator decomposition (4) becomes
(7) $e^{-\Delta\tau H}=\sum_{\vec{x}}P(\vec{x})B(\vec{x})\approx\sum_{\vec{x}:\widetilde{P}^{(n)}_{k}(\vec{x})>0}P(\vec{x})B(\vec{x})=\sum_{\vec{x}:\widetilde{P}^{(n)}_{k}(\vec{x})>0}\widetilde{P}^{(n)}_{k}(\vec{x})\left(\frac{P(\vec{x})}{\widetilde{P}^{(n)}_{k}(\vec{x})}B(\vec{x})\right).$
In a special case where $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for all $\vec{x}$, the approximation in (7) becomes exact. Plugging (7) into (5) gives
(8) $e^{-\Delta\tau H}\widehat{\Psi}^{(n)}$$\approx$ $\sum_{k}\sum_{\vec{x}:\widetilde{P}^{(n)}_{k}(\vec{x})>0}\widetilde{P}^{(n)}_{k}(\vec{x})w^{(n)}_{k}\frac{P(\vec{x})}{\widetilde{P}^{(n)}_{k}(\vec{x})}\frac{B(\vec{x}){\phi}_{k}^{(n)}}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}$
(9) $\approx$ $\sum_{k}w^{(n)}_{k}\frac{P(\vec{x}^{(n)}_{k})}{\widetilde{P}^{(n)}_{k}(\vec{x}^{(n)}_{k})}\frac{\phi_{k}^{(n+1)}}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}$
(10) $=$ $\sum_{k}\left(w^{(n)}_{k}\frac{P(\vec{x}^{(n)}_{k})}{\widetilde{P}^{(n)}_{k}(\vec{x}^{(n)}_{k})}\frac{O_{\mathrm{tr}}(\phi_{k}^{(n+1)})}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}\right)\frac{\phi_{k}^{(n+1)}}{O_{\mathrm{tr}}(\phi_{k}^{(n+1)})}$
(11) $=$ $\sum_{k}w_{k}^{(n+1)}\frac{\phi_{k}^{(n+1)}}{O_{\mathrm{tr}}(\phi_{k}^{(n+1)})},$
where
$\phi_{k}^{(n+1)}=B(\vec{x}^{(n)}_{k}){\phi}_{k}^{(n)},\vec{x}^{(n)}_{k}\sim\widetilde{P}^{(n)}_{k},$
and
(12) $w^{(n+1)}_{k}=w^{(n)}_{k}\mathcal{N}_{k}^{(n)}.$
Here, we remark that the new weight $w_{k}^{(n+1)}$ as defined in (12) does not depend on the auxiliary field $\vec{x}$ being sampled, since the normalization factor $\mathcal{N}_{k}^{(n)}$ is independent of $\vec{x}$.
We summarize the algorithm in Algorithm 1.
Algorithm 1 cp-AFQMC
a Hamiltonian $H$, initial walkers $\widehat{\Psi}^{(0)}=\sum_{k}w_{k}^{(0)}\frac{\phi_{k}^{(0)}}{O_{\mathrm{tr}}(\phi_{k}^{(0)})}$, trial wavefunction $\Psi_{\mathrm{tr}}$, step size $\Delta\tau$, number of iteration steps $M$.
Walkers $\widehat{\Psi}^{(M)}=\sum_{k}w_{k}^{(M)}\frac{\phi_{k}^{(M)}}{O_{\mathrm{tr}}(\phi_{k}^{(M)})}$.
for $n=0,1,\cdots,M-1$ do
  • For walkers satisfying $\braket{\Psi_{\mathrm{tr}},\phi^{(n)}_{k}}>0$:
    1. 1.
      Sample $\vec{x}^{(n)}_{k}\sim\frac{1}{\mathcal{N}_{k}^{(n)}}\frac{O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}P(\vec{x})$, where $B(\vec{x}^{(n)}_{k})$ is defined by the decomposition $\exp(-\Delta\tau H)=\sum_{\vec{x}}P(\vec{x})B(\vec{x})$.
    2. 2.
      $\phi_{k}^{(n+1)}=B(\vec{x}^{(n)}_{k})\phi_{k}^{(n)}$.
    3. 3.
      $w_{k}^{(n+1)}=w_{k}^{(n)}\mathcal{N}_{k}^{(n)}$.
  • Periodically compute energy as in Section 2.1.1.
  • Periodically do population control as in Section 2.1.2.
end for
In Appendix A.1, we detail the exact implementations for the transverse-field Ising (TFI) model.
2.1.1. Computing ground-state energy
For estimating the ground-state energy, we use the mixed energy estimator
$E_{\text{mixed}}=\frac{\braket{\Psi_{\mathrm{tr}},H\widehat{\Psi}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\widehat{\Psi}^{(n)}}}=\frac{\sum_{k}w^{(n)}_{k}E_{\text{local}}[\Psi_{\mathrm{tr}},\phi_{k}^{(n)}]}{\sum_{k}w^{(n)}_{k}},$
with $\widehat{\Psi}^{(n)}$ in the form of (5), and $E_{\text{local}}$ represents the local energy of a walker:
$E_{\text{local}}[\Psi_{\mathrm{tr}},\phi]=\frac{\braket{\Psi_{\mathrm{tr}},H\phi}}{\braket{\Psi_{\mathrm{tr}},\phi}}.$
2.1.2. Population control
As the number of iterations increases, some walkers may accumulate very large weights while some may have very small weights. Those walkers with small weights contribute little to the wavefunction and energy estimation, thus it is meaningless to keep track of them. To make the algorithm more efficient, we periodically implement population control [41] to adjust the weights without changing the statistical distribution. The detailed implementation is explained in Algorithm 5.
2.2. Tensor-train sketching
In this section, we introduce TT representation and TT-sketching. We consider a $d$-dimensional tensor
$u:[n_{1}]\times\cdots\times[n_{d}]\to\mathbb{R},$
where we denote $[n]=\{1,2,\cdots,n\}$ for some positive integer $n$. A TT representation is of the form
$u(i_{1},\cdots,i_{d})=\sum_{\alpha_{1}=1}^{r_{1}}\cdots\sum_{\alpha_{d-1}}^{r_{d-1}}G_{1}(i_{1},\alpha_{1})G_{2}(\alpha_{1},i_{2},\alpha_{2})\cdots G_{d-1}(\alpha_{d-2},i_{d-1},\alpha_{d-1})G_{d}(\alpha_{d-1},i_{d}),$
where we call tensor $G_{j}$ the $j$-th tensor core of the TT, and $(r_{1},\cdots,r_{d-1})$ the TT-rank. In short, we denote it as
$u=G_{1}\circ\cdots\circ G_{d}.$
We use the notation $m:n$ to denote the set $\{m,m+1,\cdots,n\}$ for two postive integers $m<n$, and further denote the truncated TT as
$G_{1:j}(i_{1:j},\alpha_{j})=\sum_{\alpha_{1}=1}^{r_{1}}\cdots\sum_{\alpha_{j-1}=1}^{r_{j-1}}G_{1}(i_{1},\alpha_{1})G_{2}(\alpha_{1},i_{2},\alpha_{2})\cdots G_{j}(\alpha_{j-1},i_{j},\alpha_{j}),$
and
$G_{j:d}(i_{j:d},\alpha_{j-1})=\sum_{\alpha_{j}=1}^{r_{j}}\cdots\sum_{\alpha_{d-1}}^{r_{d-1}}G_{j}(\alpha_{j-1},i_{j},\alpha_{j})\cdots G_{d-1}(\alpha_{d-2},i_{d-1},\alpha_{d-1})G_{d}(\alpha_{d-1},i_{d}).$
A special case is when $r_{1}=\cdots=r_{d-1}=1$, we can write the TT as
$u(i_{1},\cdots,i_{d})=G_{1}(i_{1})\cdots G_{d}(i_{d}),$
which is a separable wavefunction.
The advantage of TT is that it represents a $d$-dimensional tensor with a storage complexity linear in $d$. Moreover, the computational cost of the inner product of two TTs is also linear in $d$, which is numerically verified in Figure 1.
TT-sketching is a procedure to fit some given data into a low-rank TT. In this paper, the data type we consider is a sum of some random walkers as in (5). Each walker is a separable state. A parallel method for this task is given [38], where a left sketching function $S=S_{1}\circ\cdots\circ S_{d}$ and a right sketching function $T=T_{1}\circ\cdots\circ T_{d}$ are used to compress the data. More specifically, denote
(13) $X_{j}(\xi_{j-1},\gamma_{j-1})=\sum_{i_{1:j-1}}\sum_{i_{j:d}}S_{1:j-1}(i_{1:j-1},\xi_{j-1})\widehat{\Psi}(i_{1:j-1},i_{j:d})T_{j:d}(i_{j},\gamma_{j-1}),$
and
$Y_{j}(\xi_{j-1},i_{j},\gamma_{j})=\sum_{i_{1:j-1}}\sum_{i_{j+1}:d}S_{1:j-1}(i_{1:j-1},\xi_{j-1})\widehat{\Psi}(i_{1:j-1},i_{j},i_{j+1:d})T_{j+1:d}(i_{j+1:d},\gamma_{j}).$
The tensor cores are then given by solving the equations
$\sum_{\gamma_{j-1}}X_{j}(\xi_{j-1},\gamma_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})=Y_{j}(\xi_{j-1},i_{j},\gamma_{j}).$
This results in a high-rank TT with tensor cores $\widetilde{G}_{j}$. In practice, however, the matrices $X_{j}$ are usually low-rank. In this situation, we first apply a singular value thresholding to $X_{j}$, and then solve for the truncated tensor cores $G_{j}$, which amounts to trimming the TT to a lower rank $(r_{1},\ldots,r_{d-1})$.
For the sketching functions $S$ and $T$, we set them both to be TTs, with each tensor core $S_{j}$ and $T_{j}$ constructed as
(14) $S_{j}(\xi_{j-1},i_{j},\xi_{j})=\sum_{l}C_{S,j}(\xi_{j-1},l,\xi_{j})v_{l}(i_{j}),\ T_{j}(\xi_{j-1},i_{j},\xi_{j})=\sum_{l}C_{T,j}(\xi_{j-1},l,\xi_{j})v_{l}(i_{j}),$
where the coefficients $C_{S,j}(\xi_{j-1},l,\xi_{j})$ and $C_{T,j}(\xi_{j-1},l,\xi_{j})$ are Gaussian random tensors, and $\{b_{l}\}$ is a set of basis in the form
(15) $v_{1}=\left(1,1\right)^{T},\ v_{2}=\left(\delta,-\delta\right)^{T}.$
We aim to pick a relative small $\delta$ for the basis in (15), such that $\delta^{2}$ is still non-negligible but higher orders of $\delta$ can be ignored, then $S$ and $T$ can be considered as $1$-cluster basis sketching functions [42, 43, 44].
In addition, instead of (13), we form the matrices $X_{j}$ sequentially by making use of the tensor cores we have already obtained, of which the detailed implementation is summarized in Algorithm 2. We remark that in this algorithm, as we sequentially update the tensor cores, the computation for forming $X_{j}$ can be reused, hence the sketching algorithm has a linear time complexity in terms of the dimensionality $d$, which is also verified in Figure 1.
Figure 1. Computational cost of calculating the inner product of a TT trial wavefunction and all the walkers (upper) and TT-sketching (lower). Left: different number of spins with $2000$ walkers; right: different number of walkers for $16$ spins.
Algorithm 2 TT-sketching
A wavefunction $\widehat{\Psi}(i_{1},\cdots,i_{d})$ in the form of (5); Left sketching function $S=S_{1}\circ\cdots\circ S_{d}$, right sketching function $T=T_{1}\circ\cdots\circ T_{d}$, with $S_{j}$ and $T_{j}$ constructed as in (14).
a rank-$r$ TT that approximates $\widehat{\Psi}$ with tensor cores being $G_{1},\cdots,G_{d}$.
First phase: construct an initial TT with tensor cores $\bm{\widetilde{G}_{1},\cdots,\widetilde{G}_{d}}$
  • For the first core $j=1$:
    1. 1.
      Form $Y_{1}$:
      $Y_{1}(i_{1},\gamma_{1})=\sum_{i_{2:d}}\widehat{\Psi}(i_{1};i_{2:d})T_{2:d}(i_{2:d},\gamma_{1});$
    2. 2.
      Solve for $\widetilde{G}_{1}$:
      $\widetilde{G}_{1}(i_{1},\gamma_{1})=Y_{1}(i_{1},\gamma_{1});$
  • For the cores $j=2,\cdots,d-1$:
    1. 1.
      Form $Y_{j}$:
      $Y_{j}(\xi_{j-1},i_{j},\gamma_{j})=\sum_{i_{1:j-1}}\sum_{i_{j+1:d}}S_{1:j-1}(i_{1:j-1},\xi_{j-1})\widehat{\Psi}(i_{1:j-1},i_{j},i_{j+1:d})T_{j+1:d}(i_{j+1:d},\gamma_{j});$
    2. 2.
      Form $X_{j}$:
      $X_{2}(\xi_{1},\gamma_{1})=\sum_{i_{1}}S_{1}(x_{1},\xi_{1})\widetilde{G}_{1}(i_{1},\gamma_{1}),\ j=2;$
      $X_{j}(\xi_{j-1},\gamma_{j-1})=\sum_{\xi_{j-2}}\sum_{\gamma_{j-2}}\sum_{i_{j-1}}X_{j-1}(\xi_{j-2},\gamma_{j-2})S_{j-1}(\xi_{j-2},x_{j-1},\xi_{j-1})\widetilde{G}_{j-1}(\gamma_{j-2},i_{j-1},\gamma_{j-1}),\ j>2;$
    3. 3.
      Solve for $\widetilde{G}_{j}$:
      $\sum_{\gamma_{j-1}}X_{j}(\xi_{j-1},\gamma_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})=Y_{j}(\xi_{j-1},i_{j},\gamma_{j});$
  • For the last core $j=d$:
    1. 1.
      Form $Y_{d}$:
      $Y_{d}(\xi_{d-1},i_{d})=\sum_{i_{1:d-1}}S_{1:d-1}(i_{1:d-1},\xi_{d-1})\widehat{\Psi}(i_{1:d-1},i_{d});$
    2. 2.
      Form $X_{d}$:
      $X_{d}(\xi_{d-1},\gamma_{d-1})=\sum_{\xi_{d-2}}\sum_{\gamma_{d-2}}\sum_{i_{d-1}}X_{d-1}(\xi_{d-2},\gamma_{d-2})S_{d-1}(\xi_{d-2},x_{d-1},\xi_{d-1})\widetilde{G}_{d-1}(\gamma_{d-2},i_{d-1},\gamma_{d-1});$
    3. 3.
      Solve for $\widetilde{G}_{d}$:
      $\sum_{\gamma_{d-1}}X_{d}(\xi_{d-1},\gamma_{d-1})\widetilde{G}_{d}(\gamma_{d-1},i_{d})=Y_{d}(\xi_{d-1},i_{d});$
Second phase: compress the initial TT to targeted rank
  1. 1.
    Apply truncated SVD on $X_{j}$ by taking its $r_{j-1}$ largest singular values:
    $X_{j}(i_{j-1},\gamma_{j-1})\approx\sum_{\alpha_{j-1}=1}^{r_{j-1}}U_{j-1}(\xi_{j-1},\alpha_{j-1})\Sigma_{j-1}(\alpha_{j-1},\alpha_{j-1})V_{j-1}(\alpha_{j-1},\xi_{j-1});$
  2. 2.
    Trim the tensor cores $\widetilde{G}_{1},\cdots,\widetilde{G}_{d}$ to $G_{1},\cdots,G_{d}$ with smaller size:
    $G_{j}(\alpha_{j-1},i_{j},\alpha_{j})=\sum_{\gamma_{j-1}}\sum_{\gamma_{j}}V_{j-1}(\alpha_{j-1},\xi_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})V_{j}(\alpha_{j},\xi_{j}),2\leq j\leq d-1;$
    $G_{1}(i_{1},\alpha_{1})=\sum_{\gamma_{1}}\widetilde{G}_{1}(i_{1},\alpha_{1})V_{1}(\gamma_{1},\alpha_{1}),G_{d}(\alpha_{d-1},i_{d})=\sum_{\gamma_{d-1}}V_{d-1}(\alpha_{d-1},\xi_{d-1})\widetilde{G}_{d}(\gamma_{d-1},i_{d});$
3. Proposed Algorithm
The constraint $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}>0$ may result in a systematic bias depending on the quality of the trial function $\Psi_{\mathrm{tr}}$. Even in the case where there is no bias, having a better trial wavefunction can improve the variance. To have a better trial wavefunction, we add a step to the cp-AFQMC procedure, where we periodically update the trial function with TT-sketching using the information from the walkers in cp-AFQMC. The procedure is illustrated in Figure 2.
Formally, we summarize our proposed algorithm with details in Algorithm 3.
Figure 2. Proposed framwork. Square nodes represent random walkers, and circle nodes represent the trial wavefunction in a TT form.
Algorithm 3 cp-AFQMC with re-anchoring
A Hamiltonian $H$, trial wavefunction $\Psi_{\mathrm{tr},0}$, step size $\Delta\tau$, number of iterations $M_{\text{out}}$ and $M_{\text{in}}$, sketching functions $S$ and $T$.
An ensemble of walkers $\widehat{\Psi}_{M_{\text{out}}}=\sum_{k}w_{M_{\text{out}},k}\frac{\phi_{M_{\text{out}},k}}{O_{\mathrm{tr}}(\phi_{M_{\text{out}},k})}$, a wavefunction $\Psi_{\mathrm{tr},M_{\text{out}}}$.
Let $\widehat{\Psi}_{0}=\sum_{k}\Psi_{\mathrm{tr}}$ (i.e. $w_{0,k}=1,\phi_{0,k}=\Psi_{\mathrm{tr}}$ which implies $O_{\mathrm{tr}}(\phi_{0,k})=1$).
for $n_{\text{out}}=0,1,\cdots,M_{\text{out}}-1$ do
  • $\widehat{\Psi}_{n_{\text{out}}+1}$ = cp-AFQMC($H,\widehat{\Psi}_{n_{\text{out}}},\Psi_{\mathrm{tr},n_{\text{out}}},\Delta\tau,M_{\text{in}}$).
  • $\Psi_{\mathrm{tr},n_{\text{out}}+1}$ = TT-sketch($\widehat{\Psi}_{n_{\text{out}}+1}$,$S$,$T$).
end for
4. Convergence Analysis of cp-AFQMC
In this section, we provide a convergence analysis for cp-AFQMC to show how having a good trial wavefunction can improve the energy convergence. In Section 4.1, we first understand the effect of importance sampling in controlling the variance of the overlap of the walkers with the trial wavefunction in one step of cp-AFQMC. In Section 4.2, we then use this to show that a better trial wavefunction can lead to smaller energy variance. We further show that in cp-AFQMC, we can only hope for the convergence of energy estimator, but not the convergence of wavefunction.
4.1. Variance and bias per step of cp-AFQMC
In this section, we provide an explanation on how importance sampling controls the variance of the overlaps in one step of cp-AFQMC, and discuss the systematic bias caused by constrained-path approximation. We first consider a different representation of the walkers for the convenience of the analysis, where we define the following to be the walker that we keep track of:
(16) $\Phi_{k}^{(n)}=w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{O_{\mathrm{tr}}(\phi^{(n)}_{k})}.$
Under this notation, based on (8) (line 2), the update rule of the walker in cp-AFQMC algorithm becomes
$\Phi_{k}^{(n+1)}=\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)},$
where
(17) $\widetilde{B}_{k}^{(n)}(\vec{x}):=\frac{P(\vec{x})}{\widetilde{P}_{k}^{(n)}(\vec{x})}B(\vec{x}),\quad\vec{x}\sim\widetilde{P}_{k}^{(n)}.$
The probability distribution of importance sampling is defined in (6). It is assumed that $\widetilde{B}_{k}^{(n)}(\vec{x})$ is only well-defined when $\widetilde{P}_{k}^{(n)}(\vec{x})>0$. Again based on (8), we hope
$\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}\approx e^{-\Delta\tau H}\Phi_{k}^{(n)}$
in a certain sense. We first state a simple observation regarding the update $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$:
Lemma 1.
$\langle\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}\rangle$ is a constant in $\vec{x}$.
Proof.
Suppose we have a current walker $\Phi_{k}^{(n)}$ in the form of (16), the overlap of the walker at the $(n+1)$-th step is then
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\left\langle\Psi_{\mathrm{tr}},\frac{P(\vec{x})}{\widetilde{P}_{k}^{(n)}(\vec{x})}B(\vec{x})\Phi_{k}^{(n)}\right\rangle=\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}\mathcal{N}_{k}^{(n)}.$
The last equality is due to (6), which shows that $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ has no $\vec{x}$ dependency since $\mathcal{N}_{k}^{(n)}$ has no $\vec{x}$ dependency. ∎
Lemma 1 tells us that the variance of $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is reduced by importance sampling, in fact, to 00. We then study the value of $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$, as a consequence of Lemma 1, in the following two lemmas.
Lemma 2.
When there is no sign problem, i.e. $\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}>0$ for all $\vec{x}$, then
(18) $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}.$
Proof.
When $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}>0$, then $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for all $\vec{x}$ (see (6)). In this case, $\widetilde{B}_{k}^{(n)}(\vec{x})$ is well-defined for every $\vec{x}$ as the definition of $\widetilde{B}_{k}^{(n)}(\vec{x})$ involves division by $\widetilde{P}_{k}^{(n)}(\vec{x})$. Then,
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}}\widetilde{P}_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}}P(\vec{x})B(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}},$
where the first equality is due to Lemma 1 and the fact that $\widetilde{B}_{k}^{(n)}(\vec{x})$ can be defined for every $\vec{x}$, and the second and third equality follows from (6) and (7) respectively. ∎
It might be the case where $\widetilde{P}_{k}^{(n)}(\vec{x})=0$ for some auxiliary fields $\vec{x}$. In this situation, some of the $\vec{x}$ values cannot be reached. Hence a systematic bias on $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is introduced, as explained in the following.
Lemma 3.
When there is a sign problem, i.e. $\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\leq 0$ for some $\vec{x}$, then
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}-\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}=-\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\geq 0.$
Proof.
When $\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\leq 0$ for some $\vec{x}$, there are values of $\vec{x}$ where $\widetilde{P}_{k}^{(n)}(\vec{x})=0$. In this case, (7) can be expressed more explicitly as
$\begin{split}e^{-\Delta\tau H}&=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})>0}P(\vec{x})B(\vec{x})\\ &=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})>0}\widetilde{P}_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x}).\end{split}$
Then
$\begin{split}&\quad\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})\Phi_{k}^{(n)}+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(x)>0}\widetilde{P}_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})>0}\widetilde{P}_{k}^{(n)}(\vec{x})\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}+\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\end{split}$
where the last equality follows from Lemma 1. ∎
In this case, the actual inner product $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is larger than $\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}$, thus a systematic bias towards the trial function $\Psi_{\mathrm{tr}}$ is introduced, which causes the sign problem.
4.2. Convergence analysis
In this section, we present how importance sampling with a trial wavefunction facilitates the convergence of energy. We focus on the case where there is no sign problem as in Lemma 2. We want to show that in such a benign situation, having a trial wavefunction that is closer to the ground truth wavefunction gives a smaller variance on the energy estimator (Section 4.2.1). We further show in Section 4.2.2 that unlike typical power method, we cannot hope for the walkers to converge to the ground-state wavefunction as the variance of the walkers is too large.
The results give us a better understanding of cp-AFQMC. Although it originates from the imaginary-time evolution (1), the wavefunction itself, represented as (5), does not necessarily converge to the ground-state wavefunction. However, we can still extract useful information out of it, for example, the ground-state energy.
Suppose the Hamiltonian $H$ has eigenvalues $E_{0}<E_{1}\leq E_{2}\leq\cdots\leq E_{2^{d}-1}$, where we have assumed that the eigenspace of the smallest eigenvalue $E_{0}$ is non-degenerate. This guarantees that there is a unique lowest energy eigenvector. In this cae, the eigen-decomposition of $H$ is denoted as
$H=E_{0}\Psi_{0}\Psi_{0}^{T}+E_{1}\Psi_{1}\Psi_{1}^{T}+\cdots+E_{2^{d}-1}\Psi_{2^{d}-1}\Psi_{2^{d}-1}^{T},$
Here $\Psi_{0}$ is the ground state, and $\Psi_{i}$’s are excited states. The trial function can be decomposed using the eigenstates as
(19) $\Psi_{\mathrm{tr}}=c_{0}\Psi_{0}+c_{\perp}\Psi_{\perp},$
of which the second term $c_{\perp}\Psi_{\perp}$ represents the component of excited states, with $\Psi_{\perp}$ being a unit vector orthogonal to $\Psi_{0}$. The quantity $\left|c_{\perp}/c_{0}\right|$ then characterizes how close $\Psi_{\mathrm{tr}}$ is to the ground state $\Psi_{0}$. In what follows, we study how the convergence of the energy and the walkers depending on $\left|c_{\perp}/c_{0}\right|$.
4.2.1. Convergence analysis of the energy
In this section, we analyze how the mixed energy estimator converges to the ground-state energy in one step of cp-AFQMC. We want to show that the variance of the cp-AFQMC is controlled by $\left|c_{\perp}/c_{0}\right|$. When $\left|c_{\perp}/c_{0}\right|$ gets smaller, the variance of the energy gets smaller, i.e. the convergence of cp-AFQMC behaves like a vanilla power-iteration. In an ideal case where $\left|c_{\perp}/c_{0}\right|=0$, i.e. the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is a zero-variance principle
We first state a lemma that characterizes the propagation of a walker when applying the exact operator $e^{-\Delta\tau H}$ instead of $\widetilde{B}_{k}^{(n)}(\vec{x})$:
Lemma 4.
We further decompose the trial wavefunction as
$\Psi_{\mathrm{tr}}=c_{0}\Psi_{0}+c_{1}\Psi_{1}+\cdots+c_{2^{d}-1}\Psi_{2^{d}-1},$
and decompose the walker as
(20) $\Phi_{k}^{(n)}=c_{0}^{\prime}\Psi_{0}+c_{1}^{\prime}\Psi_{1}+\cdots+c_{2^{d}-1}^{\prime}\Psi_{2^{d}-1}.$
Assume $c_{0}c_{0}^{\prime}\neq 0$. In addition, we assume that the trial wavefunction satisfies:
(21) $a_{0}:=\left|\frac{c_{\perp}}{c_{0}}\right|=\frac{\sqrt{\sum_{j=1}^{2^{d}-1}c_{j}^{2}}}{\left|c_{0}\right|}<1,$
and
(22) $a_{1}:=\frac{\sqrt{\sum_{j=2}^{2^{d}-1}c_{j}^{2}(E_{j}-E_{0})^{2}}}{c_{1}(E_{1}-E_{0})}<1.$
Furthermore, for the walker $\Phi_{k}^{(n)}$, we denote
(23) $\tan\theta_{k}^{(n)}:=\frac{\sqrt{\sum_{j=1}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{|c^{\prime}_{0}|},\quad\tan\beta_{k}^{(n)}:=\frac{\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{|c^{\prime}_{1}|}.$
Then,
(24) $\left|\frac{\braket{\Psi_{\mathrm{tr}},He^{-\Delta\tau H}\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}-E_{0}\right|\leq e^{-\Delta\tau(E_{1}-E_{0})}\frac{1+a_{0}\tan\theta_{k}^{(n)}}{1-a_{0}\tan\theta_{k}^{(n)}}\frac{1+a_{1}\tan\beta_{k}^{(n)}}{1-a_{1}\tan\beta_{k}^{(n)}}\left|\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}-E_{0}\right|$
Proof.
The energy can be written explicitly as
$\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}=\frac{\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}E_{j}}{\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}},$
and thus
(25) $\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}-E_{0}=\frac{\sum_{j=1}^{2^{d}-1}c_{j}c_{j}^{\prime}(E_{j}-E_{0})}{\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}}.$
Similarly, we have
(26) $\frac{\braket{\Psi_{\mathrm{tr}},He^{-\Delta\tau H}\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}-E_{0}=\frac{\sum_{j=1}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}(E_{j}-E_{0})}{\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}},$
since (20) directly gives the expansion
$e^{-\Delta\tau H}\Phi_{k}^{(n)}=c_{0}^{\prime}e^{-\Delta\tau E_{0}}\Psi_{0}+c_{1}^{\prime}e^{-\Delta\tau E_{1}}\Psi_{1}+\cdots+c_{2^{d}-1}^{\prime}e^{-\Delta\tau E_{2^{d}-1}}\Psi_{2^{d}-1}.$
Now, we compare (25) and (26):
(27) $\frac{\left|\frac{\braket{\Psi_{\mathrm{tr}},He^{-\Delta\tau H}\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}-E_{0}\right|}{\left|\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}-E_{0}\right|}=\frac{\left|\sum_{j=1}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}(E_{j}-E_{0})\right|}{\left|\sum_{j=1}^{2^{d}-1}c_{j}c_{j}^{\prime}(E_{j}-E_{0})\right|}\cdot\frac{\left|\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}\right|}{\left|\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}\right|}.$
For the first term of the RHS of (27), the assumption (22) gives
(28) $\begin{split}\frac{\left|\sum_{j=1}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}(E_{j}-E_{0})\right|}{\left|\sum_{j=1}^{2^{d}-1}c_{j}c_{j}^{\prime}(E_{j}-E_{0})\right|}\leq&\frac{\left|c_{1}c_{1}^{\prime}e^{-\Delta\tau E_{1}}(E_{1}-E_{0})\right|+\left|\sum_{j=2}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}(E_{j}-E_{0})\right|}{\left|c_{1}c_{1}^{\prime}\right|(E_{1}-E_{0})-\left|\sum_{j=2}^{2^{d}-1}c_{j}c_{j}^{\prime}(E_{j}-E_{0})\right|}\\ \leq&\frac{e^{-\Delta\tau E_{1}}\left|c_{1}c_{1}^{\prime}\right|(E_{1}-E_{0})+a_{1}e^{-\Delta\tau E_{1}}|c_{1}|(E_{1}-E_{0})\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{\left|c_{1}c_{1}^{\prime}\right|(E_{1}-E_{0})-\sum_{j=2}^{2^{d}-1}\left|c_{j}c_{j}^{\prime}\right|(E_{j}-E_{0})}\\ \leq&\frac{e^{-\Delta\tau E_{1}}\left|c_{1}c_{1}^{\prime}\right|(E_{1}-E_{0})+a_{1}e^{-\Delta\tau E_{1}}|c_{1}|(E_{1}-E_{0})\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{\left|c_{1}c_{1}^{\prime}\right|(E_{1}-E_{0})-a_{1}|c_{1}|(E_{1}-E_{0})\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}\\ \leq&e^{-\Delta\tau E_{1}}\frac{1+a_{1}\frac{\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{|c^{\prime}_{1}|}}{1-a_{1}\frac{\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{|c^{\prime}_{1}|}}.\end{split}$
For the second term of the RHS of (27), we have
(29) $\begin{split}\frac{\left|\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}\right|}{\left|\sum_{j=0}^{2^{d}-1}c_{j}c_{j}^{\prime}e^{-\Delta\tau E_{j}}\right|}\leq&\frac{\left|c_{0}c_{0}^{\prime}\right|+\sum_{j=1}^{2^{d}-1}\left|c_{j}c_{j}^{\prime}\right|}{e^{-\Delta\tau E_{0}}\left|c_{0}c_{0}^{\prime}\right|-\sum_{j=1}^{2^{d}-1}\left|c_{j}c_{j}^{\prime}\right|e^{-\Delta\tau E_{j}}}\\ \leq&\frac{\left|c_{0}c_{0}^{\prime}\right|+\sum_{j=1}^{2^{d}-1}\left|c_{j}c_{j}^{\prime}\right|}{e^{-\Delta\tau E_{0}}\left|c_{0}c_{0}^{\prime}\right|-a_{0}e^{-\Delta\tau E_{0}}\left|c_{0}\right|\sqrt{\sum_{j=2}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}\\ \leq&e^{\Delta\tau E_{0}}\frac{1+a_{0}\frac{\sqrt{\sum_{j=1}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{|c^{\prime}_{0}|}}{1-a_{0}\frac{\sqrt{\sum_{j=1}^{2^{d}-1}{c_{j}^{\prime}}^{2}}}{|c^{\prime}_{0}|}}.\end{split}$
Plugging (28) and (29) back into (27) then gives (24).
Theorem 5.
Suppose there is no sign problem, as defined in Lemma 2. With the same assumptions as Lemma 4,
(30) $\left|\frac{\braket{\Psi_{\mathrm{tr}},H\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}}-E_{0}\right|=e^{-\Delta\tau(E_{1}-E_{0})}\frac{1+a_{0}\tan\theta_{k}^{(n)}}{1-a_{0}\tan\theta_{k}^{(n)}}\frac{1+a_{1}\tan\beta_{k}^{(n)}}{1-a_{1}\tan\beta_{k}^{(n)}}\left|\frac{\braket{\Psi_{\mathrm{tr}},H\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}-E_{0}\right|\\ +\|H-E_{0}I\|_{2}\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\frac{a_{0}\tan\theta_{k}^{(n)}}{1+a_{0}e^{-\Delta(E_{1}-E_{0})}\tan\theta_{k}^{(n)}}.$
Proof.
We split the energy into a deterministic part and a fluctuation part:
(31) $\left|\frac{\braket{\Psi_{\mathrm{tr}},H\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}}-E_{0}\right|\leq\left|\frac{\braket{\Psi_{\mathrm{tr}},He^{-\Delta\tau H}\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}-E_{0}\right|+\left|\frac{\braket{\Psi_{\mathrm{tr}},H(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}\right|,$
where we have used (18) for the denominator. The bound on the first term directly follows from Lemma 4. For the second term, plugging in the decomposition (19) gives
(32) $\frac{\braket{\Psi_{\mathrm{tr}},H(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(33) $=$ $\frac{\braket{H\Psi_{\mathrm{tr}},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(34) $=$ $\frac{\braket{c_{0}E_{0}\Psi_{0}+c_{\perp}H\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(35) $=$ $\frac{\braket{c_{\perp}(H-E_{0}I)\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(36) $=$ $\frac{\braket{c_{\perp}(H-E_{0}I)\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(37) $\leq$ $c_{\perp}\|H-E_{0}I\|_{2}\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}|}$
(38) $=$ $c_{\perp}\|H-E_{0}I\|_{2}\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|c_{0}\braket{\Psi_{0},e^{-\Delta\tau H}\Phi_{k}^{(n)}}+c_{\perp}\braket{\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}}|}$
(39) $\leq$ $\|H-E_{0}I\|_{2}\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\left|\frac{c_{\perp}}{c_{0}}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{\braket{\Psi_{0},\Phi_{k}^{(n)}}}\right|\bigg{/}\left(1+a_{0}e^{-\Delta(E_{1}-E_{0})}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$
(40) $=$ $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\frac{a_{0}\tan\theta_{k}^{(n)}\|H-E_{0}I\|_{2}}{1+a_{0}e^{-\Delta(E_{1}-E_{0})}\tan\theta_{k}^{(n)}},$
where second equality follows from (21) and Cauchy–Schwarz inequality, and we use $\mathcal{P}_{\perp}$ to denote the orthogonal projector that projects to the space spanned by $\{\Psi_{1},\ldots\Psi_{2^{d}-1}\}$. Plugging this and the bound in Lemma 4 into (31) then gives (30). ∎
We notice that the energy convergence consists of a deterministic power iteration part and a fluctuating part that is rougly $a_{0}\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}=\left|c_{\perp}/c_{0}\right|\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}$ (when $a_{0}$ is small). Here $\|H-E_{0}I\|_{2}$ resembles the energy scale, and $\tan\theta_{k}^{(n)}$ is a random variable that captures the deviation of a random walker from the ground state. While currently, we do not have a characterization of the variance of $\tan\theta_{k}^{(n)}$, if we assume that it has a bounded variance of $O(1)$, when $a_{0}$ gets small, i.e. the trial function gets closer to the ground-state wavefunction, the fluctuation term can be made small. In an ideal case where $a_{0}=0$ so that the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is no fluctuation in the energy estimation. As a matter of fact, we have
$\frac{\braket{\Psi_{0},H\Psi}}{\braket{\Psi_{0},\Psi}}=E_{0},$
for any wavefunction $\Psi$.
4.2.2. Convergence analysis of the walkers
Now, we analyze the convergence of a particular walker $\Phi_{k}^{(n)}$ to the ground state, by keeping track of the angle between $\Phi_{k}^{(n)}$ and the ground state $\Psi_{0}$. As shown later in Theorem 6, unlike the results in energy convergence, even if the trial function is exactly the ground-state wavefunction, there is still an error term that does not vanish. This implies that a single walker may not be able to converge to somewhere near the ground state, even with a good trial wavefunction.
We denote the tangent of the angle between $\Phi_{k}^{(n)}$ and $\Psi_{0}$ as $\tan\theta_{k}^{(n)}$ as in (23). Similarly, the tangent of the angle between $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$ ($\widetilde{B}_{k}^{(n)}(\vec{x})$ is defined in (17)) and $\Psi_{0}$ is denoted as $\tan\theta_{k}^{(n+1)}(\vec{x})$. Our goal is to show that there is no contraction in terms of this angle in 1-step of cp-AFQMC. To make the notations more concise, we rewrite the angle defined in (23) as
$\tan\theta_{k}^{(n+1)}(\vec{x})=\frac{\|\mathcal{P}_{\perp}\widetilde{B}_{k}^{(n)}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}|},\quad\tan\theta_{k}^{(n)}=\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|},$
where $\mathcal{P}_{\perp}$ is the orthogonal projector that projects to the space spanned by $\{\Psi_{1},\ldots\Psi_{2^{d}-1}\}$. We want to show in the following that $\tan\theta_{k}^{(n+1)}(\vec{x})<\tan\theta_{k}^{(n)}$ up to some noise terms. However, the noise terms will not vanish even for very good trial wavefunction.
Theorem 6.
Suppose there is no sign problem, as defined in Lemma 2. With the same assumptions as Lemma 4, we further assume that
$a_{0}<e^{-\Delta\tau E_{0}}\frac{|{\braket{\Psi_{0},\Phi_{k}^{(n)}}}|}{|{\braket{\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}|}},$
Then,
$\tan\theta_{k}^{(n+1)}(\vec{x})\leq e^{-\Delta\tau(E_{1}-E_{0})}\tan\theta_{k}^{(n)}+e^{\Delta\tau E_{0}}\left(a_{0}\tan\theta_{k}^{(n)}+1\right)O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)\\ \quad+a_{0}e^{2\Delta\tau E_{0}}O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)^{2},$
Proof.
To estimate $\tan\theta_{k}^{(n+1)}(\vec{x})$, we need to calculate each component of $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$ under the basis $\{\Psi_{j}\}_{j}$. We first study the overlap of $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$ with the ground state $\Psi_{0}$
(41) $\braket{\Psi_{0},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\\ \left\langle\frac{\Psi_{\mathrm{tr}}-c_{\perp}\Psi_{\perp}}{c_{0}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}\right\rangle=\frac{1}{c_{0}}\langle\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}\rangle-\frac{c_{\perp}}{c_{0}}\braket{\Psi_{\perp},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}.$
For the first term, we have
$\begin{split}&\quad\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=\braket{c_{0}\Psi_{0}+c_{\perp}\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=c_{0}\braket{e^{-\Delta\tau E_{0}}\Psi_{0},\Phi_{k}^{(n)}}+c_{\perp}\braket{\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}},\\ \end{split}$
where in the first equality we use Lemma 2, in the second equality we use (19), and in the third equality we use the fact that $e^{-\Delta\tau H}$ is a Hermitian matrix. Plugging this back into (41) we get
$\begin{split}\braket{\Psi_{0},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=e^{-\Delta\tau E_{0}}\braket{\Psi_{0},\Phi_{k}^{(n)}}-\frac{c_{\perp}}{c_{0}}\braket{\Psi_{\perp},E^{(n)}_{k}}.\end{split}$
Now, the component of $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$ in the excited subspace can be characterized as
$\begin{split}\|\mathcal{P}_{\perp}\widetilde{B}_{k}^{(n)}\Phi_{k}^{(n)}\|&\leq\|\mathcal{P}_{\perp}e^{-\Delta\tau H}\Phi_{k}^{(n)}\|+\|\mathcal{P}_{\perp}\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|\leq e^{-\Delta\tau E_{1}}\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|+\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|,\\ \end{split}$
where $\mathcal{P}_{\perp}$ projects to
$\begin{split}\tan\theta_{k}^{(n+1)}(\vec{x})&=\frac{\|\mathcal{P}_{\perp}\widetilde{B}_{k}^{(n)}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}|}\\ &\leq\frac{e^{-\Delta\tau E_{1}}\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|+\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{e^{-\Delta\tau E_{0}}|\braket{\Psi_{0},\Phi_{k}^{(n)}}|-\left|\frac{c_{\perp}}{c_{0}}\right||\braket{\Psi_{\perp},\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}}|}\\ &\leq e^{-\Delta\tau(E_{1}-E_{0})}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\left(1+e^{\Delta\tau E_{0}}\left|\frac{c_{\perp}}{c_{0}}\right|O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)\right)\\ &\quad+\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{e^{-\Delta\tau E_{0}}\braket{\Psi_{0},\Phi_{k}^{(n)}}}\left(1+e^{\Delta\tau E_{0}}\left|\frac{c_{\perp}}{c_{0}}\right|O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)\right)\\ &\leq e^{-\Delta\tau(E_{1}-E_{0})}\tan\theta_{k}^{(n)}+e^{\Delta\tau E_{0}}\left(\tan\theta_{k}^{(n)}\left|\frac{c_{\perp}}{c_{0}}\right|+1\right)O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)\\ &\quad+e^{2\Delta\tau E_{0}}\left|\frac{c_{\perp}}{c_{0}}\right|O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)^{2}.\end{split}$
As we can see, even when $a_{0}=0$, i.e. the trial wavefunction is exactly the ground-state wavefunction, there is still an error term $O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$. Thus, the walker is not guaranteed to get closer to the ground state. In fact, generally, one needs exponential number of walkers to get $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}<1$. This is due to the fact that generally, one needs to sample an exponential number of $\widetilde{B}_{k}^{(n)}(\vec{x})$ to approximate $\exp(-\Delta\tau H)$.
5. Numerical Results
In this section, we apply our proposed algorithm to both 1D and 2D Transverse-Field Ising (TFI) systems. For more details about TFI system, we refer readers to Appendix A. For the 1D system, we consider a periodic boundary condition, while for the 2D system, we use an open boundary condition.
In all of our experiments, the step size of the imaginary-time evolution is set to be $\Delta\tau=0.01$. We set the initial trial wavefunction to be the fully disordered state
(42) $\Psi_{\mathrm{tr}}=\frac{1}{\sqrt{2^{d}}}\cdot\left(1,1\right)^{T}\otimes\cdots\otimes\left(1,1\right)^{T},$
which can also be interpreted as the eigenvector corresponding to the smallest eigenvalue of $H_{1}$ (The first term of the Hamiltonian in (43)). We implement TT-sketching to update the trial wavefunction every $50$ steps. At the $2000$-th step, we stop updating the trial wavefunction and keep using the latest one. We set the number of walkers to scale linearly with the number of spins. For systems with $16,32,64$ and $96$ spins, the numbers of walkers are set to be $2000,4000,8000$ and $12000$, respectively.
In the TT-sketching procedure, we take $\delta=0.1$ in (15) for the sketching functions. For systems with $16$ or $32$ spins, the rank of the sketching functions is set to be $60$. For systems with $64$ or $96$ spins, we set the rank to be $150$. The target rank of the sketched TT is set to be $r_{1}=r_{d-1}=2$, and $r_{2}=\cdots=r_{d-2}=4$ in all experiments.
The 1D transverse-field Ising model with periodic boundary undergoes a quantum phase transition [45] at $g=1$ (see (43)), which is demonstrated in Figure 3 using the 1D system with $16$ spins as an example. To be more specific, We run our algorithm on the 1D system with $16$ spins with different transverse fields, and measure the ground-state energy $E_{0}(g)$. We also calculate the energy of the transverse field by
$\frac{\braket{\Psi_{0}(g),(-\sum_{i=1}^{d}\sigma_{i}^{x})\Psi_{0}(g)}}{\braket{\Psi_{0}(g),\Psi_{0}(g)}}=\frac{\mathrm{d}}{\mathrm{d}g}E_{0}(g),$
where $\Psi_{0}(g)$ represents the ground-state wavefunction of the system with the transverse-field strength being $g$. Numerically, this derivative is approximated by finite difference:
$\frac{\mathrm{d}}{\mathrm{d}g}E_{0}(g)\approx\frac{E_{0}(g+\Delta g)-E_{0}(g)}{\Delta g},$
where we pick $\Delta g=0.01$ in practice.
Figure 3. Results for $16$ spins in 1D with different transverse-field strength $g$. Left: ground-state energy; right: the energy of the transverse field, where the critical point $g=1$ is illustrated with the black dashed line.
Since the quantum phase transition at the critical point $g=1$ makes it the most difficult case to deal with, we also compare cp-AFQMC with re-anchoring and the vanilla cp-AFQMC on larger systems, including 1D system with $d=32,64$ and $96$ spins at $g=1$. The ground-state energy of 1D periodic system can be calculated analytically [46], and the ground-state wavefunction can be well approximated by density matrix renormalization group (DMRG) [47, 48].
In addition, we also test a quasi-1D system, where the spins are placed on a $4\times 16$ lattice. We incorporate periodic boundary on the $4$-site axis, and open boundary on the $16$-site axis. Thus, it appears like a cylinder. The transverse field is also set to be $g=1$. Despite that the spins are arranged on a 2D lattice, the ground-state energy and ground-state wavefunction can still be well-approximated by DMRG, due to the specific type of boundary conditions. Thus, we still use DMRG result as a reference.
As illustrated in Figure 4, for each system, the energy obtained by vanilla cp-AFQMC is of the order $O(10^{-3})$, while cp-AFQMC with re-anchoring can achieve the accuracy of order $O(10^{-5})$ or $O(10^{-6})$. The relative error of the energy is shown in Table 1. Besides, the overlap between our optimized trial wavefunction and the ground-state wavefunction is shown in Figure 5. For 1D systems, the optimized trial wavefunction in the form of a rank-$4$ TT can approximate the ground-state wavefunction quite well. For systems with $32$, $64$ and $96$ spins, the overlap can achieve around $0.9$, $0.8$, and $0.7$ respectively. For the $4\times 16$ cylinder system, the trial wavefunction may converge to somewhere not quite close to the ground-state wavefunction. However, we can still get pretty accurate ground-state energy using the optimized trial wavefunction.
Figure 4. Comparison between the energy convergence of cp-AFQMC and cp-AFQMC with re-anchoring at $g=1$. Top: cp-AFQMC; bottom: cp-AFQMC with re-anchoring. From left to right: $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
Figure 5. The overlap between our optimized trial wavefunction and the ground-state wavefunction. From left to right: $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
cp-AFQMC cp-AFQMC with re-anchoring
$32$ spins 1D $(+1.35\pm 0.31)\times 10^{-3}$ $(+0.44\pm 2.43)\times 10^{-5}$
$64$ spins 1D $(+1.88\pm 0.37)\times 10^{-3}$ $(+0.77\pm 1.07)\times 10^{-5}$
$96$ spins 1D $(+2.49\pm 0.33)\times 10^{-3}$ $(-4.95\pm 8.94)\times 10^{-6}$
$4\times 16$ spins 2D (cylinder) $(+4.18\pm 1.42)\times 10^{-3}$ $(-0.66\pm 1.96)\times 10^{-6}$
Table 1. Relative error of the energy.
From the results in Table 1, we can tell that there is an obvious bias in the estimation of the ground-state energy by vanilla cp-AFQMC. As discussed in detail in Appendix B, the bias originates from Trotter error and population control.
Finally, we test our algorithm on the $11\times 1$ 11 lattice with open boundary at the quantum critical point $g\approx 3.044$ [49]. In our experiments, we take $g=3$. The results are shown in Figure 6. For such a 2D system, there is no analytic solution for the ground-state energy anymore. A benchmark for the energy density given by DMRG is $-3.17210\pm 1\times 10^{-5}$ [50], while the energy density estimated by our method is $-3.17212\pm 4\times 10^{-5}$.
Figure 6. Energy convergence of cp-AFQMC with re-anchoring for $11\times 1$ 11 lattice.
6. Conclusion
In this paper, we proposed a framework for calculating the ground-state energy of quantum many-body systems. Our algorithm improved cp-AFQMC by periodically updating the trial wavefunction using TT-sketching. As indicated by numerical results, our algorithm is accurate for the transverse-field Ising model. Furthermore, we provided a convergence analysis for cp-AFQMC for the first time. A potential work in the future is to apply this framework to electronic systems.
Appendix A Transverse-Field Ising Model
We take the transverse-field Ising (TFI) model [51, 52] as an example. We follow the procedure of the CPMC algorithm on Hubbard model [41] to solve for TFI model. The Hamiltonian of TFI model is given by
(43) $H=-g\sum_{i=1}^{d}\sigma_{i}^{x}-\sum_{\langle i,j\rangle}\sigma_{i}^{z}\sigma_{j}^{z}:=H_{1}+H_{2},$
where $g$ is the strength of the magnetic field, $d$ is the total number of spins, $\langle i,j\rangle$ represents pairs of nearest neighbors, $\sigma^{x},\sigma^{z}\in\mathbb{R}^{2\times 2}$ are Pauli matrices [53], and we denote
$\sigma_{i}^{x,z}=I_{2}\otimes\cdots\otimes\underbrace{\sigma^{x,z}}_{i\text{-th term}}\otimes\cdots\otimes I_{2}\in\mathbb{R}^{2^{d}\times 2^{d}}.$
For such a Hamiltonian, we use the Suzuki-Trotter approximation [54, 55] to approximate $e^{-\Delta\tau H}$:
(44) $e^{-\Delta\tau H}=e^{-\Delta\tau H_{1}/2}e^{-\Delta\tau H_{2}}e^{-\Delta\tau H_{1}/2}+O(\Delta\tau^{3}):=B_{H_{1}/2}B_{H_{2}}B_{H_{1}/2}+O(\Delta\tau^{3}),$
where we refer to $B_{H_{1}/2}$ and $B_{H_{2}}$ as the one-body and two-body propagator, respectively. For the one-body propagator, it holds
$B_{H_{1}/2}=e^{g\Delta\tau\sigma^{x}/2}\otimes e^{g\Delta\tau\sigma^{x}/2}\otimes\cdots\otimes e^{g\Delta\tau\sigma^{x}/2}.$
Now we need to decompose the two-body propagator $B_{H_{2}}$ in order to have the decomposition in (4). The trick is to use a discrete Hubbard-Stratonovich transformation [56]:
(45) $e^{\Delta\tau\sigma_{i}^{z}\sigma_{j}^{z}}=e^{-\Delta\tau}\frac{1}{2}\sum_{x_{ij}=\pm 1}e^{x_{ij}\lambda_{ij}(\sigma_{i}^{z}+\sigma_{j}^{z})},$
where the constants $\lambda_{ij}$ are given by $\cosh{2\lambda_{ij}}=e^{2\Delta t}$, and each bond $\langle i,j\rangle$ is corresponded with an auxiliary field $x_{ij}$. We rewrite (45) as
$e^{\Delta\tau\sigma_{i}^{z}\sigma_{j}^{z}}=\sum_{x_{ij}=\pm 1}p(x_{ij})b(x_{ij}),$
where we treat $x_{ij}$ as a random variable and $p(x_{ij})=1/2$ for $x_{ij}=\pm 1$ as its probability distribution function, and
$b(x_{ij})=e^{-\Delta\tau}e^{x_{ij}\lambda_{ij}(\sigma_{i}^{z}+\sigma_{j}^{z})}=e^{-\Delta\tau}e^{x_{ij}\lambda_{ij}\sigma_{i}^{z}}e^{x_{ij}\lambda_{ij}\sigma_{j}^{z}}$
is a product of two separable operators, thus it returns another separable state when acting on one. The two-body propagator can then be rewritten as
(46) $B_{H_{2}}=e^{\Delta\tau\sum_{\braket{i,j}}\sigma_{i}^{z}\sigma_{j}^{z}}=\prod_{\braket{i,j}}e^{\Delta t\sigma_{i}^{z}\sigma_{j}^{z}}=\prod_{\braket{i,j}}\sum_{x_{ij}=\pm 1}p(x_{ij})b(x_{ij}),$
which yields the desired decomposition (4). Here, the matrices $\sigma_{i}^{z}\sigma_{j}^{z}$ are diagonal, thus the equality is exact in (46), and no extra Trotter error will be introduced. Since the Hamiltonian contains two different terms $B_{H_{2}}$ and $B_{H_{1}/2}$, to make cp-AFQMC more efficient, we modify Algorithm 1 slightly as follows.
Algorithm 4 cp-AFQMC-TFI
Hamiltonians $H_{1}$ and $H_{2}$; initial walkers $\widehat{\Psi}^{(0)}=\sum_{k}w_{k}^{(0)}\frac{\phi_{k}^{(0)}}{O_{\mathrm{tr}}(\phi_{k}^{(0)})}$; trial wavefunction $\Psi_{\mathrm{tr}}$, step size $\Delta\tau$, number of iteration steps $M$.
Walkers $\widehat{\Psi}^{(M)}=\sum_{k}w_{k}^{(M)}\frac{\phi_{k}^{(M)}}{O_{\mathrm{tr}}(\phi_{k}^{(M)})}$
for $n=0,1,\cdots,M-1$ do
  • For walkers with positive weights:
    1. 1.
      Let $\phi^{(n)\prime}_{k}\leftarrow B_{H_{1}/2}\phi^{(n)}_{k}$, $w_{k}^{(n)}\leftarrow w_{k}^{(n)}\frac{O_{\mathrm{tr}}(\phi^{(n)\prime}_{k})}{O_{\mathrm{tr}}(\phi^{(n)}_{k})}$.
    2. 2.
      Update $\phi_{k}^{(n)\prime}$ to $\phi_{k}^{(n)\prime\prime}$ by applying $B_{H_{2}}$ to $\phi_{k}^{(n)\prime}$ stochastically using the decomposition in (46). Update weights $w_{k}^{(n)}$ to $w_{k}^{(n+1)}$. The details are provided in Section A.1.
    3. 3.
      Let $\phi^{(n+1)}_{k}\leftarrow B_{H_{1}/2}\phi^{(n)\prime\prime}_{k}$, $w_{k}^{(n+1)}\leftarrow w_{k}^{(n+1)}\frac{O_{\mathrm{tr}}(\phi^{(n+1)}_{k})}{O_{\mathrm{tr}}(\phi^{(n)\prime\prime}_{k})}$.
  • Periodically compute energy as in Section 2.1.1.
  • Periodically do population control as in Section 2.1.2.
end for
We remark that all of the entries in both the matrices $B_{H_{1}/2}$ and $B_{H_{2}}$ are non-negative. Thus, if we apply $B_{H_{1}/2}B_{H_{2}}B_{H_{1}/2}$ to a wavefunction with all the entries being positive, it returns another all-positive-entry function.
A.1. Details for applying $B_{H_{2}}$
In this section, we detail how to apply $B_{H_{2}}$ to a walker $\phi_{k}^{(n)\prime}$ stochastically. To summarize, we sample $x_{ij}$ for each $\langle i,j\rangle$ one at a time, and apply a single $b(x_{ij})$ to $\phi^{(n)\prime}_{k}$. For example, for a single pair $\langle i,j\rangle$, we let
$\sum_{x_{ij}}p(x_{ij})b(x_{ij})\frac{w_{k}^{(n)}\phi_{k}^{(n)\prime}}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})}=\sum_{x_{ij}}\widetilde{p}_{ij,k}^{(n)}(x_{ij})w_{k}^{(n)}\frac{p(x_{ij})}{\widetilde{p}_{ij,k}^{(n)}(x_{ij})}\frac{b(x_{ij})\phi_{k}^{(n)\prime}}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})}\\ \approx w_{k}^{(n)}\frac{p(x^{(n)}_{ij,k})}{\widetilde{p}_{ij,k}^{(n)}(x^{(n)}_{ij,k})}\frac{b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime}}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})},$
where we obtain $x^{(n)}_{ij,k}$ by importance sampling according to a new distribution $\widetilde{p}_{ij,k}^{(n)}$. Similar to Section 2.1, we let
$\widetilde{p}_{ij,k}^{(n)}(x_{ij})=\frac{1}{\mathcal{N}^{(n)}_{ij,k}}\frac{O_{\mathrm{tr}}(b(x_{ij})\phi_{k}^{(n)\prime})}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})}p(x_{ij}),$
where $p(x_{ij})=1/2$ for $x_{ij}=\pm 1$. In this case, the update rule for the weight and the walker becomes
$w^{(n)}_{k}\leftarrow\begin{cases}w_{k}^{(n)}\mathcal{N}^{(n)}_{ij,k}\quad&\text{if}\ \langle\Psi_{\mathrm{tr}},b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime}\rangle>0,\\ 0&\text{otherwise},\end{cases}$
and
$\phi^{(n)\prime\prime}_{k}=b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime}.$
This procedure is then repeated for all pairs $\langle i,j\rangle$.
Appendix B Systematic bias of vanilla cp-AFQMC on TFI system
In this section, we study the origin of the bias in this section, using the TFI system with $16$ spins in 1D as an example.
The good news is that for TFI system, with a properly chosen trial wavefunction $\Psi_{\mathrm{tr}}$, the cp-AFQMC is free of sign problem. In fact, as shown in Appendix A, the operator $B(\vec{x})$ in (4) has the property that if it acts on a wavefunction $\phi$ with all entries being positive, the result $B(\vec{x})\phi$ still has all positive entries. Therefore, by choosing a trial function with all positive entries, the walkers will remain such a form, thus it always holds $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}>0$. The trial wavefunction (42) satisfies the condition, thus there will be no sign problem if one runs cp-AFQMC with this fixed trial function.
Without sign problem, the constrained-path approximation will not lead to any systematic bias by zeroing some of the walkers. However, there are other issues that will bring a bias to the energy estimation. One of them is the Trotter error as shown in (44). At each imaginary-time step, the Trotter error is $O(\Delta\tau^{3})$. For a fixed imaginary-time, one needs to run $O(1/\Delta\tau)$ steps of cp-AFQMC. Thus, the total error in terms of the energy estimation accumulates to $O(\Delta\tau^{2})$.
Ideally, the Trotter error will vanish as the imaginary-time step $\Delta\tau$ goes to 00. However, this is not always the case as shown in Figure 7(a). With $2000$ walkers, the estimated energy indeed converges to the ground-state energy as $\Delta\tau$ goes to 00. If we use $200$ walkers, however, there is still a quite obvious bias even when $\Delta\tau$ becomes 00. Moreover, with fewer walkers, the bias is consistently larger with any imaginary-time step $\Delta\tau$.
(a) Trotter error
(b) Finite-population error
Figure 7. Results of cp-AFQMC on $16$-spin system in 1D. (a): estimated energy with different imaginary-time step $\Delta\tau$; (b): estimated energy with different population size.
To further study how the error depends on the number of walkers, we fix the imaginary-time step to be $\Delta\tau=0.01$, and run cp-AFQMC with different population size. As shown in Figure 7(b), the error is roughly inversely proportional to the population size. The bias is introduced by the population control procedure summarized in Algorithm 5. To be more specific, if we look at each single walker, it breaks into a random number of copies, with the average number of copies matching its weight. However, to maintain a constant population size, the number of copies must be adjusted based on the total, which introduces dependency among walkers. This dependency leads to a bias in the energy estimation.
Algorithm 5 Population control
walkers $\phi_{k}$, weights $w_{k}$, and $O_{\mathrm{tr}}(\phi_{k})$.
adjusted walkers $\phi_{k}$, weights $w_{k}$, and $O_{\mathrm{tr}}(\phi_{k})$.
normalize the weights so that the average value of the weights is $1$.
$s\leftarrow\text{a uniformly random number in }(0,1)$,
$n_{1}\leftarrow 1$,
for $k=1,2,\cdots,N$ do
     $s\leftarrow s+w_{k}$;
     $n_{2}\leftarrow\text{the smallest integer that is greater than or equal to }s$;
     set the walkers $\phi_{n_{1}},\cdots,\phi_{n_{2}-1}$ to be $\phi_{k}$, and corresponding overlaps to be $O_{\mathrm{tr}}(\phi_{k})$;
     $n_{1}\leftarrow n_{2}$
end for
reset all the weights to be $w_{k}=1$.